Grade 12Mechanics

a plate of mass M length and breadth a is initially lying on a horizontal floor with the line parallel to the floor and the breadth perpendicular to the floor and find the work done to erect it on its breadth?

Brijendra
8 Years agoGrade 12

1 Answer

Aastha Mishra
6 Years ago
Gravitational potential energy of the block at position 1= mgh1
At position 2=mgh2
WD by external agent=W=∆U=mgh2-mgh1
Where h2=b/2  and h1=a/2 as it is a regular shaped object
So W = mg(b/2-a/2)